# Zero-Shot Machine-Generated Text Detection Using Mixture of Large Language Models

### v1

@@ Abstract
The dissemination of Large Language Models (LLMs), trained at scale, and endowed with powerful text-generating abilities has vastly increased the threats posed by generative AI technologies by reducing the cost of producing harmful, toxic, faked or forged content. In response, various proposals have been made to automatically discriminate artificially generated from human-written texts, typically framing the problem as a classification problem. Most approaches evaluate an input document by a well-chosen detector LLM, assuming that low-perplexity scores reliably signal machine-made content. As using one single detector can induce brittleness of performance, we instead consider several and derive a new, theoretically grounded approach to combine their respective strengths. Our experiments, using a variety of generator LLMs, suggest that our method effectively increases the robustness of detection.
@@ 1 Introduction
Large Language Models (LLMs) have greatly improved the fluency and diversity of machine-generated texts. The release of ChatGPT and GPT-4 by OpenAI has sparked global discussions on the effective use of AI-based writing assistants. However, this progress has also introduced considerable threats such as fake news Zellers et al. (2019), and the potential for harmful outputs such as toxic or dishonest speech (Crothers et al., 2023), among others. As it seems, the research on methods aimed at detecting the origin of a given text to mitigate the dissemination of forged content and to prevent technology-aided plagiarism still lags behind the rapid advancement of AI itself. [^1] As illustrated by the discontinuation of OpenAI’s own detector https://openai.com/index/new-ai-classifier-for-indicating-ai-written-text/.
@@
Many works have focused on tools that could spot such AI-generated outputs and address these underlying risks. From a bird’s eye view, this typically involves using detector models to discriminate generator models’ outputs from legitimate human writings. Multiple versions of this generic text classification task have been considered, varying e.g. the number of possible outcomes and the amount of available supervision (see Section 5). Owing to its large user base and applications, the largest effort has focused on one specific generator, ChatGPT, for which training and test data is easily obtained. Yet, the corresponding supervised binary problem, with a unique known generator, is not the only way to frame this task. A more challenging problem, that we study here, is generator-agnostic artificial text detection, where the models to detect are not known in advance.
@@
As pointed out e.g. in Antoun et al. (2024); Hans et al. (2024); Wang et al. (2024a), the performance of artificial text detection systems varies depending on the choices of the detector(s) / generator(s) pair. This means, in principle, that the search for optimal detection performance should include a systematic exploration of the space of possible detectors. As the number and diversity of LLMs keep increasing, such exploration seems not only challenging but also unrealistic. Furthermore, Dugan et al. (2024) demonstrated shown that the current detection methods are easily fooled when changing the generator or the sampling method. Our proposal thus considers ensemble methods, where a coalition of several available models can be exploited to build the detector. This approach eliminates the need to empirically search for the best detector(s), and yields detection systems that can robustly detect multiple generators. Furthermore, it lends itself to the smooth extension of the ensemble as new models become available while improving generalization to unseen models.
@@
Our main contribution. In this paper, using fundamental information-theoretic principles from universal compression, we derive a new ensemble method that optimally combines the strength of multiple LLMs . Relying on existing benchmarks, as well as a new corpus of machine-generated texts, we confirm in our experiments that this method yields detectors that can robustly detect a multiplicity of generators and that it compares favorably with several recent proposals. Our analyses explore the effect of adding models into the ensemble, and also highlight, using a multilingual dataset, how the contribution of each constituent model changes when the source of artificial texts is modified.
2 Robust Scoring Algorithms (RSA)
@@ 2.1 Background
We consider models for language generation tasks that define a probability distribution over strings. Formally, these language models are probability distributions over an output space $\mathcal{Y}$ which contains all possible strings over vocabulary $\Omega$: $\mathcal{Y}\triangleq\big{\{}\texttt{BOS}\circ\mathbf{y}\circ\texttt{EOS}\,|\,\mathbf{y}\in\Omega^{*}\big{\}},$ BOS and EOS denote for special reserved beginning-of-sequence and end-of-sequence tokens, respectively, and $\Omega^{*}$ denotes the Kleene closure of $\Omega$.
@@
Measurement of information. The fundamental concept in information is “surprisal”, using the relationship: $\textrm{information}=-\log(\textrm{probability})$  Cover and Thomas (2006), and assuming the use of coding techniques such as Huffman and Arithmetic codes Shields (1996) which give message lengths closely approximating the ideal length in binary digits. Here, the measure of information is most conveniently introduced in the context of lossless compression. That is, we will look a what happens when information is passed from a encoder to a decoder. Humans often use codes, such as natural languages, which are not optimal for any set of prior expectations. There are good reasons for natural languages to be less than “optimal” when regarded as codes for data compression. One reason is that spoken language is transmitted from speaker to listener via a noisy channel. Codes with substantial redundancy can tolerate some degree of corruption without becoming unintelligible, whereas optimal compression codes, in which every digit matters, are very sensitive to corruption. Artificial codes for information storage are often designed so that legal strings conform to a strict pattern while most strings do not. A corrupted received string can then be corrected by replacing it with the nearest legal string, provided the degree of corruption is not too great. However, for our purposes, we need not consider errors in the storage of messages, and hence need not be concerned with error-correcting codes.
@@
Finally, another important concept is the conditional mutual information (MI) between two random variables $\mathbb{M}$ and $Y_{t}$, given a sequence value $\mathbf{y}_{<t}$, defined as Cover and Thomas (2006):
@@
2.2 Combining LLMs with Robust Scoring Algorithms (RSA)
@@
A family of LLMs can be exploited to produce explanations of token sequences. To this end, we assume some $\hat{m}_{t}=f_{t}(\mathbf{y}_{<t})$ which selects a probability distribution $p_{\theta}(y_{t}|\hat{m}_{t},\mathbf{y}_{<t})$ over $\Omega$. Given $\hat{m}_{t}$, the encoder can construct an optimum code for token $y_{t}$, using distribution $p_{\theta}(y_{t}|\hat{m}_{t},\mathbf{y}_{<t})$. Thus, a rich family of LLMs therefore allows us to capture and represent regular patterns in token sequences via the model selector $\hat{m}_{t}$ and subsequently use it to minimize the total expected codelength.
Identifying explanations of data. We now turn to the problem of determining an adequate sequence of models $\hat{\mathbf{m}}=\langle\hat{m}_{0},\ldots,\hat{m}_{T}\rangle$. Our goal will be to derive a robust scoring algorithm that best extracts regularity in the data, which is equivalent to identifying the model that achieves the best compression of the input tokens. Suppose we are given a family of LLMs $\mathcal{P}_{\mathcal{M}}(\mathcal{Y})$, with corresponding Shannon codelengths $\mathcal{L}_{\theta}(y_{t}|m,\mathbf{y}_{<t})\triangleq-\log p_{\theta}(y_{t}|m,\mathbf{y}_{<t})$ for each $y_{t}$. These can be viewed as a collection of data compressors, indexed by $m$. We can measure the performance of encoding $y_{t}$ at time $t$ relative to $\mathcal{P}_{\mathcal{M}}(\mathcal{Y})$. If we chose to encode the token $y_{t}$ using a model $q(y_{t}|\mathbf{y}_{<t})$, the resulting expected excess codelength (or overhead) w.r.t. any distribution $p_{\theta}\in\mathcal{P}_{\mathcal{M}}(\mathcal{Y})$ is given by:
@@
Leveraging codelengths for identifying AI-generated text. The averaged overhead of the optimal codelength $-\log q^{\star}(y_{t}|\mathbf{y}_{<t})$ obtained by solving (3) seems to be a very reasonable choice for building a robust score function to detect AI-generated text because of the following properties:
@@
Definition 1 (Robust Scoring Algorithm).
The Robust Scoring Algorithm (RSA) $S_{\textrm{Av}}(\mathbf{w})$ for an input sentence $\mathbf{w}=\langle w_{0},w_{1},\ldots\rangle$ is defined as:
$S_{\textrm{Av}}(\mathbf{w})\triangleq\frac{1}{T}\displaystyle\sum_{t=1}^{T}\Big{[}\underbrace{\sum_{y_{t}\in\Omega}\mathbbm{1}[y_{t}=w_{t}]\mathcal{L}^{\star}(y_{t}|\mathbf{w}_{<t})}_{\textrm{(codelength for observed token)}}-\frac{1}{M}\sum_{m\in\mathcal{M}}\underbrace{\Big{(}\sum_{y_{t}\in\Omega}p_{\theta}(y_{t}|m,\mathbf{w}_{<t})\mathcal{L}^{\star}(y_{t}|\mathbf{w}_{<t})\Big{)}}_{\textrm{(codelength for generated token using model `m')}}\Big{]},$
@@
Remark 2.
An alternative score, denoted by ${S}_{\max}(\mathbf{w})$, can be defined by replacing the sum over $m$ with its maximum. ${S}_{\textrm{Av}}(\mathbf{w})$ and ${S}_{\max}(\mathbf{w})$ will be empirically evaluated.
The next proposition provides a theoretically result together with an efficient iterative algorithm to optimally solve expression (3). The proof of this proposition is relegated to Appendix A.
@@
The optimal solution to (3) is shown to be the mixture of LLMs:
@@
where the distribution $\mu^{\star}(\cdot|\mathbf{y}_{<t})$ of the random variable $\mathbb{M}$ of LLM index on $\mathcal{M}$ follows by solving:
@@
Furthermore, the weights $\{\mu^{\star}(m|\mathbf{y}_{<t})\}_{m\in\mathcal{M}}$ can be computed efficiently by using Blahut–Arimoto algorithm, referred to as Blahut–Arimoto weights.
@@ 2.3 Implementation
Proposition 1 implies that in order to implement the scoring function introduced in Definition 1, it would be enough to solve the optimization in (4) which is much simpler than (3). Interestingly, Blahut–Arimoto algorithm (Arimoto, 1972; Blahut, 1972) provides us with an efficient iterative method to to compute the maximization of mutual information (4) (see Appendix B). This algorithm lies at the core of our scoring procedure.
Algorithm 1 RSA Scoring
@@
7:$S(\mathbf{w})\leftarrow\frac{1}{T}\sum_{t}s_{t}(\mathbf{w})$triangleright RSA score over the whole text
@@ 3.1 Datasets & metrics
To evaluate our method on a diverse set of texts and generative models, we collected four publicly available datasets used in previous publications: Ghostbuster (Verma et al., 2023), Binoculars (Hans et al., 2024), M4 (Wang et al., 2024a) and a corpus of scholarly texts Liyanage et al. (2022).
@@
  • The Ghostbuster dataset is split into three parts: WritingPrompts, based on the r/WritingPrompts subreddit where users submit stories in response to short prompts; Reuters, using the Reuters 50-50 authorship identification dataset (Houvardas and Stamatatos, 2006); and Essays, comprised of essays scraped from IvyPanda, [^2] https://ivypanda.com/ a website dedicated to homework help. Each part contains 1,000 original texts that have been regenerated using headlines with ChatGPT (using 5 different prompts) and with Claude (Anthropic, 2023) (one prompt), for a total of 7,000 texts.
@@
  • The Binoculars dataset contains samples of human-written texts from CCNews, Pubmed and CNN; alternative completions are automatically generated using Llama-2-13b (Touvron et al., 2023) and Falcon-7b (Almazrouei et al., 2023). Their generation technique collects the first 50 tokens for each text sample and uses them as a prompt to generate machine output. Those first 50 tokens are then removed from the result so that samples only contain machine-generated texts.
@@
  • The M4333For Multi-Lingual, Multi-Domain, Multi-Generator Machine-Generated text detection. paper describes a massive dataset of natural texts collected from a diverse set of sources. Comparable artificial texts are obtained by generating with 6 LLMs, with prompts such as article titles, headlines, or abstracts depending on their domain. In our experiments, we only use the two “multilingual” generators (ChatGPT  [^4] https://chatgpt.com/ and Davinci003555https://platform.openai.com/docs/models), and the balanced sets of $3,000$ pairs of (artificial, natural) texts in Chinese, Russian, Bulgarian, Arabic, and German.
@@
  • The academic benchmark proposed by Liyanage et al. (2022) is generated using a GPT2 model Radford et al. (2019) fine-tuned on papers from Arxiv’s “Computation and Language" (CS.CL) domain. 100 seed texts are used to generate new papers of similar length, using the first 50 words as prompts. Sections such as “methodology”, “results”, “evaluation”, and “discussion” are voluntarily omitted from texts, to ensure that discriminating factors do not include the comparative use of diagrams, tables, and equations.
@@
Table 1 displays the main statistics for these 4 test sets. For completeness, we also augmented the Binoculars dataset using Llama-2-7b as an alternative text generator with Huggingface’s transformers model.generate() (Wolf et al., 2020). [^6] https://huggingface.co/docs/transformers. Sampling: repetition_penalty: 1.18, temp.: 0.6, top_p: 0.9. As for the original corpus, the first 50 tokens of the original texts provide the starting context for both machine & human texts. To test our method on extreme cases, we also generated 3000 texts of 500-token length using a unigram model trained on the Brown Corpus (Francis and Kucera, 1979), along with original extracts of the same length for comparison, details are in section  4.1.4. These new corpora will be released with our detector’s implementation.
@@
M4 (Multilingual) ChatGPT, Davinci-003. 24,000 729 24,000 649
Table 1: Human [H] and Artificial [A] texts used in the experiments, all lengths are in Llama-2 tokens.
Taken together, these datasets represent a large variety of genres, styles, themes, languages, and generators, allowing us to thoroughly assess the strengths and weaknesses of our detection strategy.
@@
As in most studies, we report the AUROC score as our main evaluation metric. Depending on the application, True Positive Rate (TPR) for a predefined False Positive rate (e.g., 5%) is also worth looking at and is also reported in most of our results. It is important to note that we compare imbalanced datasets in the case of Ghostbuster and the original Binoculars datasets, see numbers in Table  1, whereas, for our Binoculars regenerations and the M4 multilingual datasets, the data is comprised of (original, generated) pairs of texts.

### v2

# MOSAIC: Multiple Observers Spotting AI Content, a Robust Approach to Machine-Generated Text Detection

@@ Abstract
The dissemination of Large Language Models (LLMs), trained at scale, and endowed with powerful text-generating abilities has vastly increased the threats posed by generative AI technologies by reducing the cost of producing harmful, toxic, faked or forged content. In response, various proposals have been made to automatically discriminate artificially generated from human-written texts, typically framing the problem as a classification problem. Most approaches evaluate an input document by a well-chosen detector LLM, assuming that low-perplexity scores reliably signal machine-made content. As using one single detector can induce brittleness of performance, we instead consider several and derive a new, theoretically grounded approach to combine their respective strengths. Our experiments, using a variety of generator LLMs, suggest that our method effectively leads to robust detection performances. An early version of the code is available at https://github.com/BaggerOfWords/MOSAIC.
@@ 1 Introduction
Large Language Models (LLMs) have greatly improved the fluency and diversity of machine-generated texts. The release of ChatGPT and GPT-4 by OpenAI has sparked global discussions regarding the effective use of AI-based writing assistants. This progress has also introduced considerable threats such as fake news generation Zellers et al. (2019), and the potential for harmful outputs such as toxic or dishonest content (Crothers et al., 2023), among others. As it seems, the research on methods aimed at detecting the origin of a given text to mitigate the dissemination of forged content and to prevent technology-aided plagiarism still lag behind the rapid advancement of AI itself. [^1] As illustrated by the discontinuation of OpenAI’s detector https://openai.com/index/new-ai-classifier-for-indicating-ai-written-text/.
@@
Many works have focused on tools that could spot such AI-generated outputs and address these underlying risks. From a bird’s eye view, this typically involves using detector models to discriminate generator models’ outputs from legitimate human writings. Multiple versions of this generic text classification task have been considered, varying e.g. the number of possible categories to distinguish and the amount of supervision (see Section 5). Owing to its large user base and applications, the largest effort has focused on one specific generator, ChatGPT, for which training and test data is easily obtained. Yet, the corresponding supervised binary problem, with a unique known generator, is not the only way to frame this task. A more challenging problem, that we study here, is generator-agnostic artificial text detection, where the models to detect are not known in advance.
@@
As pointed out e.g., in (Antoun et al., 2024; Hans et al., 2024; Wang et al., 2024a), the performance of artificial text detection systems varies depending on the choices of the detector(s) / generator(s) pair. The detector may serve to assess text probabilities, as in (Mitchell et al., 2023; Bao et al., 2024), or to regenerate texts, as e.g., in (Mao et al., 2024; Yang et al., 2024). This implies that the search for optimal detection performance should include a systematic exploration of the space of possible detectors. As the number and diversity of LLMs keep increasing, such exploration seems not only challenging but also unrealistic. Furthermore, Dugan et al. (2024) demonstrated that the current detection methods are brittle and easily fooled by simply changing the generator or the associated sampling method, which means that the optimal detector may need to be periodically revised.
All this suggests that more efforts are needed to increase the robustness of existing detectors to changes in the generation method. For this, our proposal relies on ensemble methods, where a coalition of several models can be exploited to build the detector. For this, we generalize perplexity-based approaches, which flag as “artificial” texts that have a suspiciously small perplexity. As perplexity is also an encoding measure, our combination algorithm will seek to identify time-varying mixture models to minimize the worst-case expected encoding size. This also corresponds to the combination leading to the highest mutual information, that we implement with an architecture depicted in Figure 1. Further details, explanations, and proofs can be found in section 2. This approach eliminates the need to empirically search for the best detector(s), and yields detection systems that can robustly detect multiple generators. Furthermore, it lends itself to the smooth enrichment of the ensemble as new models become available, thereby improving generalization to unseen models.
Figure 1: Mixture Model. $\{\mu_{i}\}$ are weights associated to models in the mixture, defined in Equation 4
@@
Our contributions. In this paper, using fundamental information-theoretic principles from universal compression, we derive a new ensemble score that optimally combines the strength of multiple LLMs to detect forged texts. Our experiments use both standard benchmarks comprising multiple domains and languages, as well as a new corpus of machine-generated texts. They confirm that ensembling strong LLMs yields detectors that can robustly identify a multiplicity of generators and that compare favorably with several recent proposals using a predefined set of detector models. Our analyses explore the effect of incrementally adding models into the ensemble, and also highlight, using a multilingual dataset, how the contribution of each constituent model changes when the source of artificial texts is modified.
2 The MOSAIC approach
@@ 2.1 Background
We consider models for language generation tasks that define a probability distribution over strings. Formally, language models are probability distributions over an output space $\mathcal{Y}$ which contains all possible strings over vocabulary $\Omega$: $\mathcal{Y}\triangleq\big{\{}\texttt{BOS}\circ\mathbf{y}\circ\texttt{EOS}\,|\,\mathbf{y}\in\Omega^{*}\big{\}},$ BOS and EOS denote respectively the beginning-of-sequence and end-of-sequence tokens, and $\Omega^{*}$ is the Kleene closure of $\Omega$.
@@
Measuring information. The fundamental concept in information is “surprisal”, using the relationship: $\textrm{information}=-\log(\textrm{probability})$  (Cover and Thomas, 2006), and assuming the use of coding techniques such as Huffman and Arithmetic codes (Shields, 1996) which give message lengths closely approximating the ideal length in binary digits. Here, the measure of information is most conveniently introduced in the context of lossless compression. That is, we will look a what happens when information is passed from a encoder to a decoder. Humans often use codes, such as natural languages, which are not optimal for any set of prior expectations. There are good reasons for natural languages to be less than “optimal” when regarded as codes for data compression. One reason is that spoken language is transmitted from speaker to listener via a noisy channel. Codes with substantial redundancy can tolerate some degree of corruption without becoming unintelligible, whereas optimal compression codes, in which every digit matters, are very sensitive to corruption. Artificial codes for information storage are often designed so that legal strings conform to a strict pattern while most strings do not. A corrupted received string can then be corrected by replacing it with the nearest legal string, provided the degree of corruption is not too great. However, for our purposes, we need not consider errors in the storage of messages, nor be concerned with error-correcting codes.
@@
Finally, another important concept is the conditional mutual information (MI) between two random variables $\mathbb{M}$ and $Y_{t}$, given a sequence value $\mathbf{y}_{<t}$, defined as (Cover and Thomas, 2006):
@@
2.2 Combining LLMs
@@
A family of LLMs can be exploited to produce explanations of token sequences. To this end, we assume some $\hat{m}_{t}=f_{t}(\mathbf{y}_{<t})$ which selects a probability distribution $p_{\theta}(y_{t}|\hat{m}_{t},\mathbf{y}_{<t})$ over $\Omega$. Given $\hat{m}_{t}$, the encoder can construct an optimum code for token $y_{t}$, using distribution $p_{\theta}(y_{t}|\hat{m}_{t},\mathbf{y}_{<t})$. Therefore, a rich family of LLMs allows us to capture and represent regular patterns in token sequences via the model selector $\hat{m}_{t}$ and subsequently use it to minimize the total expected codelength.
Identifying explanations of data. We now turn to the problem of determining an adequate sequence of models $\hat{\mathbf{m}}=\langle\hat{m}_{0},\ldots,\hat{m}_{T}\rangle$. Our goal will be to derive a robust scoring algorithm that best extracts regularity in the data, which is equivalent to identifying the model that achieves the best compression of the input tokens. Suppose we are given a family of LLMs $\mathcal{P}_{\mathcal{M}}(\mathcal{Y})$, with corresponding Shannon codelengths $\mathcal{L}_{\theta}(y_{t}|m,\mathbf{y}_{<t})\triangleq-\log p_{\theta}(y_{t}|m,\mathbf{y}_{<t})$ for each $y_{t}$. These can be viewed as a collection of data compressors, indexed by $m$. We can measure the performance of encoding $y_{t}$ at time $t$ relative to $\mathcal{P}_{\mathcal{M}}(\mathcal{Y})$. If we chose to encode the token $y_{t}$ with model $q(y_{t}|\mathbf{y}_{<t})$, the resulting expected excess codelength (or overhead) w.r.t. any distribution $p_{\theta}\in\mathcal{P}_{\mathcal{M}}(\mathcal{Y})$ is:
@@
Leveraging codelengths for identifying AI-generated text. The averaged overhead of the optimal codelength $-\log q^{\star}(y_{t}|\mathbf{y}_{<t})$ obtained by solving Eq. (3) seems to be a very reasonable choice for building a robust score function to detect AI-generated text because of the following properties:
@@
Definition 1 (MOSAIC Score).
For an input sentence $\mathbf{w}=\langle w_{0},w_{1},\ldots\rangle$, the MOSAIC score is defined as:
$S_{\textrm{Av}}(\mathbf{w})\triangleq\frac{1}{TM}\displaystyle\sum_{t=1}^{T}\sum_{m\in\mathcal{M}}\Big{[}\underbrace{\sum_{y_{t}\in\Omega}\mathbbm{1}[y_{t}=w_{t}]\mathcal{L}^{\star}(y_{t}|\mathbf{w}_{<t})}_{\textrm{(codelength for observed token)}}-\underbrace{\sum_{y_{t}\in\Omega}p_{\theta}(y_{t}|m,\mathbf{w}_{<t})\mathcal{L}^{\star}(y_{t}|\mathbf{w}_{<t})}_{\textrm{(codelength for generated tokens from model `m')}}\Big{]},$
@@
The next proposition provides a theoretical result together with an efficient iterative algorithm to optimally solve expression (3). The proof of this proposition is relegated to Appendix A.
@@
The optimal solution to (3) is a mixture of LLMs:
@@
where the distribution $\mu^{\star}(\cdot|\mathbf{y}_{<t})$ of the random variable $\mathbb{M}$ over LLM indice in $\mathcal{M}$ satisfies:
@@
Furthermore, the weights $\{\mu^{\star}(m|\mathbf{y}_{<t})\}_{m\in\mathcal{M}}$ can be computed efficiently with the Blahut–Arimoto algorithm, and are referred to as Blahut–Arimoto weights.
@@ 2.3 Implementation
Proposition 1 implies that to implement the scoring function introduced in Definition 1, it would be enough to solve the optimization in Eq. (4), which is much simpler than Eq. (3). Interestingly, Blahut–Arimoto algorithm (Arimoto, 1972; Blahut, 1972) provides us with an efficient iterative method to compute the maximization of mutual information in Eq. (4) (see Appendix B). This algorithm lies at the core of our scoring procedure.
Algorithm 1 MOSAIC Scoring
@@
7:$S_{\textrm{Av}}(\mathbf{w})\leftarrow\frac{1}{T}\sum_{t}s_{t}(\mathbf{w})$triangleright MOSAIC score for the whole text
@@ 3.1 Datasets & metrics
We evaluate our method on a diverse set of texts and generative models from the literature: RAID (Dugan et al., 2024), Ghostbuster (Verma et al., 2023), Binoculars (Hans et al., 2024), M4 (Wang et al., 2024a) and a corpus of scholarly texts (Liyanage et al., 2022).
RAID contains about 15k natural texts in English from a variety of domains; the artificial part version contains approximately 500K, generated with a diverse set of recent models, also varying the sampling procedure. As the test set is not publicly released, we select a balanced random subset of 2000 texts for our experiments. RAID also includes an artificially noised subcorpus, which was not used in our experiments.
@@
The Ghostbuster dataset is split into three parts: WritingPrompts, based on the r/WritingPrompts subreddit where users submit stories in response to short prompts; Reuters, using the Reuters 50-50 authorship identification dataset (Houvardas and Stamatatos, 2006); and Essays, comprising essays scraped from IvyPanda, [^2] https://ivypanda.com/ a website dedicated to homework help. Each part contains 1,000 original texts that have been regenerated from their headlines [^3] For Reuters and Essays, the headlines were themselves generated based on the text. with ChatGPT (using 5 different prompts) and with Claude (Anthropic, 2023) (one prompt), for a total of 7,000 texts.
@@
The Binoculars dataset contains samples of human-written texts from CCNews, Pubmed and CNN; alternative completions are automatically generated using a Llama-2-13b (Touvron et al., 2023) and Falcon-7b (Almazrouei et al., 2023). Their generation technique uses the first 50 tokens of each text as a prompt to generate a machine output. Those first 50 tokens are then removed from the result so that samples only contain machine-generated texts. [^4] For most texts, only one single artificial text (using either Llama or Falcon) is available.
@@
The M4555For Multi-Lingual, Multi-Domain, Multi-Generator Machine-Generated text. corpus is a massive dataset of natural texts collected from a diverse set of sources. Comparable artificial texts are generated by 6 LLMs, with prompts such as article titles, headlines, or abstracts depending on their domain. In our experiments, we only use one “multilingual” generator (ChatGPT, https://chatgpt.com/), and the balanced sets made of $3,000$ pairs of (artificial, natural) texts in Russian (Rus), Bulgarian (Bul), Arabic (Ara), and German (Ger).
@@
The academic benchmark proposed by Liyanage et al. (2022) is generated using a GPT2 model (Radford et al., 2019) fine-tuned on papers from Arxiv’s “Computation and Language" (CS.CL) domain. 100 seed texts are used to generate new papers of comparable length, using the first 50 words as prompts. Sections such as “methodology”, “results”, “evaluation”, and “discussion” are voluntarily omitted from texts, to ensure that discriminating factors do not rely on the comparative use of diagrams, tables, and equations.
@@
Table 1 displays the main statistics for these 5 corpora. For completeness, we also augment the Binoculars dataset using Llama-2-7b and Mistral-7b as alternative text generators with Huggingface’s transformers model.generate() (Wolf et al., 2020). [^6] https://huggingface.co/docs/transformers. For Llama: nucleus sampling parameters: repetition_penalty: 1.18, temp.: 0.6, top_p: 0.9; for Mistral: repetition_penalty: 1.18, temp.: 0.7, top_p: 1. As for the original corpus, the first 50 tokens of the original texts provide the starting context for both machine & human texts. To also test our method on extreme cases, we randomly generated 3,000 texts of 500-token using a unigram model trained on the Brown Corpus (Francis and Kucera, 1979), [^7] We tokenize the Brown Corpus (Francis and Kucera, 1979) with Llama tools and use count ratios as probability estimates, with Laplace smoothing ($\epsilon=1e-10$). along with original extracts of the same length. These corpora will be released with our detector’s implementation.
Table 1: Natural and Artificial texts used in the experiments, all lengths are in Llama-2 tokens.
@@
M4 (Multilingual) ChatGPT 12,000 729 12,000 649
These datasets represent a large variety of genres, themes, languages, sampling strategies, and generators, allowing us to thoroughly assess the robustness of our detection strategy. Using Binoculars+, we can evaluate detection performance for texts produced by one of our detectors.
@@
As in most studies, we report the AUROC score as our main evaluation metric. Depending on the application, True Positive Rate (TPR) for a predefined False Positive rate (e.g., 5%) is also worth looking at and is also reported in most of our results. All these scores are obtained using scikit-learn (Pedregosa et al., 2011). It is important to note that our experiments involve imbalanced settings in the case of Ghostbuster and the original Binoculars datasets, see numbers in Table 1, whereas, for our Binoculars+ regenerations and the M4 multilingual datasets, the test data contains a balanced number of (original, generated) pairs of texts.